# Mixer is more than just a model

### 28 Feb 2024 v1

@@ Abstract
Recently, MLP structures have regained popularity, with MLP-Mixer standing out as a prominent example. In the field of computer vision, MLP-Mixer is noted for its ability to extract data information from both channel and token perspectives, effectively acting as a fusion of channel and token information. Indeed, Mixer represents a paradigm for information extraction that amalgamates channel and token information. The essence of Mixer lies in its ability to blend information from diverse perspectives, epitomizing the true concept of ”mixing” in the realm of neural network architectures. Beyond channel and token considerations, it is possible to create more tailored mixers from various perspectives to better suit specific task requirements. This study focuses on the domain of audio recognition, introducing a novel model named Audio Spectrogram Mixer with Roll-Time and Hermit FFT (ASM-RH) that incorporates insights from both time and frequency domains. Experimental results demonstrate that ASM-RH is particularly well-suited for audio data and yields promising outcomes across multiple classification tasks.
@@ III-C CASIA Chinese Sentiment Corpus
The best model was selected based on metrics from the validation dataset, and the evaluated metrics were reported for the test set. Three different random seeds were used for the validation process of each model. The evaluation metrics encompassed accuracy (ACC, primary metric) and area under the curve (AUC, secondary metric). The mean values of the experimental results are shown in Table III.
TABLE III: CASIA Result
@@ III-D RAVDESS
RAVDESS (Ryerson Audio-Visual Database of Emotional Speech and Song) is an emotional speech and song database contributed by participants from Ryerson University, Canada. This dataset includes recordings of audio clips featuring various emotional states. It comprises 24 performers, evenly split between males and females, who deliver speech and songs depicting emotions such as happiness, neutrality, sadness, anger, surprise, fear, disgust, and calmness. In this section, we employ a 10-fold cross-validation method and consider the best performance achieved across the folds as the final result.
TABLE IV: RAVDESS Result
@@ III-E Ablation Study
To assess the effectiveness of Roll-Time-mixing and Hermit-Frequency-mixing, ablation experiments were conducted on the UrbanSound8K dataset. Channel-mixing and Tokens-mixing in the MLP-Mixer were implemented as alternatives to Roll-Time-mixing and Hermit-Frequency-mixing, respectively. The experiment maintained consistency by utilizing the same three random seeds, ensuring reproducibility and reliability. The results were averaged and are presented in Table V for a comprehensive comparison and analysis. In the table, ”-H” represents ASM-RH with Hermit-Frequency-mixing only, while ”-R” signifies ASM-RH with Roll-Time-mixing only.
@@
TABLE V: UrbanSound8K Ablation Experiment Result

### 02 Mar 2024 v2

@@ Abstract
Recently, MLP structures have regained popularity, with MLP-Mixer standing out as a prominent example. In the field of computer vision, MLP-Mixer is noted for its ability to extract data information from both channel and token perspectives, effectively acting as a fusion of channel and token information. Indeed, Mixer represents a paradigm for information extraction that amalgamates channel and token information. The essence of Mixer lies in its ability to blend information from diverse perspectives, epitomizing the true concept of ”mixing” in the realm of neural network architectures. Beyond channel and token considerations, it is possible to create more tailored mixers from various perspectives to better suit specific task requirements. This study focuses on the domain of audio recognition, introducing a novel model named Audio Spectrogram Mixer with Roll-Time and Hermit FFT (ASM-RH) that incorporates insights from both time and frequency domains. Experimental results demonstrate that ASM-RH is particularly well-suited for audio data and yields promising outcomes across multiple classification tasks. The models and optimal weights files will be published.
@@ III-B UrbanSound8K
To compare with the current state-of-the-art (SOTA) methods, we performed a 10-fold cross-validation experiment on the UrbanSound 8K dataset. The average and best results are presented in Table III for reference.
TABLE III: UrbanSound8K 10-fold Result
Avg acc(%) Best acc(%)
EAT-M 90 -
ASM-RH-I 97.96 98.63
@@ III-C CASIA Chinese Sentiment Corpus
The best model was selected based on metrics from the validation dataset, and the evaluated metrics were reported for the test set. Three different random seeds were used for the validation process of each model. The evaluation metrics encompassed accuracy (ACC, primary metric) and area under the curve (AUC, secondary metric). The mean values of the experimental results are shown in Table IV.
TABLE IV: CASIA Result
@@ III-D RAVDESS
RAVDESS (Ryerson Audio-Visual Database of Emotional Speech and Song) is an emotional speech and song database contributed by participants from Ryerson University, Canada. This dataset includes recordings of audio clips featuring various emotional states. It comprises 24 performers, evenly split between males and females, who deliver speech and songs depicting emotions such as happiness, neutrality, sadness, anger, surprise, fear, disgust, and calmness. In this section, we employ a 10-fold cross-validation method and consider the best performance achieved across the folds as the final result. The result is in Table V.
TABLE V: RAVDESS Result
@@ III-E Ablation Study
To assess the effectiveness of Roll-Time-mixing and Hermit-Frequency-mixing, ablation experiments were conducted on the UrbanSound8K dataset. Channel-mixing and Tokens-mixing in the MLP-Mixer were implemented as alternatives to Roll-Time-mixing and Hermit-Frequency-mixing, respectively. The experiment maintained consistency by utilizing the same three random seeds, ensuring reproducibility and reliability. The results were averaged and are presented in Table VI for a comprehensive comparison and analysis. In the table, ”-H” represents ASM-RH with Hermit-Frequency-mixing only, while ”-R” signifies ASM-RH with Roll-Time-mixing only.
@@
TABLE VI: UrbanSound8K Ablation Experiment Result